You travel 2 hour at 35 mph, 3 hours at 55 mph and 4 hours at 70 mph. What is your average speed?
Found 3 solutions by stanbon, solver91311, palanisamy:
Answer by stanbon(75887)

You travel 2 hour at 35 mph, 3 hours at 55 mph and 4 hours at 70 mph. What is your average speed?
---
rate = (total distance)/(total time)
---
total distance = 2*35+3*55+4*70 = 515 miles
total time = 2+3+4 = 9 hrs
-----
rate = 515/9 = 57.22 mph
==============================
Cheers,
Stan H.

Answer by solver91311(24713)



If you travel 2 hours at 35 mph, then you travel 70 miles. 3 hours at 55 means you travel 165 miles, and 4 hours at 70 is 280 miles. 70 plus 165 plus 280 is 515 miles. It took 9 hours to travel 515 miles, hence average speed is 515 divided by 9 or just a touch over 57.2 mph.

John


Answer by palanisamy(496)

You travel 2 hour at 35 mph, 3 hours at 55 mph and 4 hours at 70 mph.
Total distance travelled = 2*35+3*55+4*70
= 70+165+280
= 515 miles
Total time = 2+3+4 = 9 hours
Average speed = 515/9 = 57 2/9 miles/hour